OpenStudy (anonymous):

3x + 2y = 4 2x - y = 5 This system of equations has no solution. has one solution. correct? is coincident.

OpenStudy (kamibug):

Your guess is right. It has one solution.

OpenStudy (afifarahman):

has one solution i.e (2, -1)

OpenStudy (kamibug):

Yep. If you have a graphin program and you insterted the lines' equations you;d see that that is the intersection (aka: solution)

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

3x + 2y = 4 2y = -3x + 4 y = -3/2x + 2 2x - y = 5 2x - 5 = y yes I agree. One solution